# Pose Prior Learner: Unsupervised Categorical Prior Learning for Pose Estimation

### v2

@@ 3.4 Training and Inference
Our PPL is trained to jointly minimize all the four losses: the image reconstruction loss $L_{ir}$, the boundary loss $L_{b}$, the link regularization loss $L_{l}$, and the keypoint configuration reconstruction loss $L_{kr}$. We elaborate on these four losses below. Additionally, we use three training techniques for stable convergence of the network. We present the details of these training techniques in Appendix A6.
Image Reconstruction Loss. If PPL correctly estimates the pose on the original image $I$, the reconstructed image $I_{recon}$, based on the estimated pose, should be identical to $I$. Therefore, ensuring the quality of $I_{recon}$ encourages PPL to improve its pose estimation accuracy. To achieve this, we apply a perceptual loss on the embeddings of $I$ and $I_{recon}$, extracted using a frozen feature extractor $\psi(\cdot)$ from the VGG19 network pre-trained on ImageNet (Russakovsky et al., 2015). The perceptual loss is defined as: $L_{ir}=\|\psi(I_{recon})-\psi(I)\|_{1}$.
Boundary Loss. To ensure that the network does not transform the points in the keypoint prior outside the boundaries of the image, we limit the x and y coordinates of the transformed keypoints to be within the image:
$L_{b}=\sum_{*\in{x,y}}\begin{cases}|P^{\prime}_{i,*}|&\text{if }|P^{\prime}_{i,*}|>1,\\ 0&\text{otherwise}.\end{cases}$ (5)
where $P^{\prime}_{i,x}$ and $P^{\prime}_{i,y}$ are the normalized $x$ and $y$ coordinate of the keypoint $P^{\prime}_{i}$ respectively.
Link Regularization Loss. A person’s arm always maintains a fixed length regardless of the poses. Thus, we propose the constraint that links should be assigned a high weight if they do not vary significantly in length before and after the affine transformation. The loss $L_{l}$ encourages the preservation of the link lengths during pose estimation. It is defined as in the equation below, where $l(\cdot)$ is the $L2$ distance between two keypoints before and after the affine transformation:
$L_{l}=\sum_{i,j}w_{i,j}\|l(P_{i},P_{j})-l(P^{\prime}_{i},P^{\prime}_{j})\|_{1}.$ (6)
@@
Table 1: Keypoint detection on CUB-200-2011. We report the mean $L2$ error normalized by the image resolutions of $128\times 128$. We use 10 keypoints for CUB-Aligned and 4 keypoints for CUB-001, CUB-002, CUB-003, and CUB-all. The best is in bold.
Method CUB- aligned CUB- 001 CUB- 002 CUB- 003 CUB- all
(Zhang et al., 2018) 5.36 26.9 27.6 27.1 22.4
(He et al., 2021) 5.21 22.6 29.1 21.2 14.7
(He et al., 2022b) 3.23 22.1 22.3 21.5 12.1
AutoLink (He et al., 2022a) 3.51 20.2 19.2 18.5 11.3
PPL (ours) 3.19 19.3 18.6 17.3 10.5
Table 2: Keypoint detection on Human3.6m. We report the mean $L2$ error normalized by the image resolutions (Res.) of both $256\times 256$ and $128\times 128$. * is the result trained by using masked images as reference images. The best is in bold.
Method Res. Norm. $L2$ Error
(Jakab et al., 2020) 256 2.73
(Thewlis et al., 2017) 256 7.51
(Lorenz et al., 2019) 256 2.79
(Zhang et al., 2018) 256 4.91
(Schmidtke et al., 2021) 256 3.31
AutoLink (He et al., 2022a) 128 2.76
PPL (ours) 128 1.92
PPL* (ours) 128 2.23
PPL (ours) 256 2.56
Table 3: Keypoint detection on Taichi. For consistency with the baseline methods, we report the summed $L2$ error at a resolution of $256\times 256$. * is the result trained by using masked images as reference images. The best is in bold.
Method Summed $L2$ Error
(Siarohin et al., 2021) 389.78
(He et al., 2021) 437.69
(Zhang et al., 2018) 343.67
(He et al., 2022b) 417.17
AutoLink (He et al., 2022a) 316.10
PPL* (ours) 298.60
PPL (ours) 293.35
Iterative Inference. We propose an iterative inference strategy (Figure 3). $4$ iterations are used for every experiment. In every iteration, we take the reconstructed image $I_{recon}$ from the last iteration (the original image $I$ for iteration 00) as the input. We infer its keypoints $T^{\prime}$ as the output keypoints of the current iteration. The hierarchical memory $M$ is used to reconstruct $T^{\prime}$ and the reconstructed keypoints $T^{\prime}_{recon}$ are used to obtain the reconstructed image $I_{recon}$. $I_{recon}$ is then used as the input for the next iteration. We keep the original occluded image $I$ as the reference image for all the iterations. See Appendix A1 for more implementation details.
Figure 3: Overview of the iterative inference strategy in our PPL (Section 3.4). During inference, we iteratively use the reconstructed image $I_{recon}$ as input to estimate the pose $T^{\prime}$. The hierarchical memory $M$ refines the estimated pose $T^{\prime}$ and outputs $T^{\prime}_{recon}$. With the original image $I$ as the reference image $I_{ref}$, PPL reconstructs the image $I_{recon}$. It is then used as the input image in the next iteration.
@@ 4 Experiments
For quantitative experiments, we use three image datasets: Human3.6m (Ionescu et al., 2013), Taichi (Siarohin et al., 2019b, a), and CUB-200-2011 (Wah et al., 2011). For fair comparisons, we set the number of keypoints $N$ to be the same as (He et al., 2022a). For qualitative visualizations, we use Youtube dog videos, Flowers (Nilsback & Zisserman, 2008), 11k-Hands (Afifi, 2019), and Horses (Zhu et al., 2017). See Appendix A2 for details. We utilize video frames as $I_{ref}$ for the Human3.6m, Taichi, and YouTube dog videos, while randomly masked images are used as $I_{ref}$ for other datasets. On Human3.6m and CUB-200-2011, we report the results in the mean $L2$ error between the predicted keypoints and the ground truth, normalized by the image size. For Taichi (He et al., 2022a; Siarohin et al., 2021; He et al., 2021, 2022b; Zhang et al., 2018), we use the summed $L2$ error computed at a resolution of $256\times 256$.
@@ 4.1 Unsupervised Categorical Pose Estimation
We compare the keypoint detection results of PPL with other unsupervised pose estimation methods and present the results in Table 1 (CUB-200-2011), Table 2 (Human3.6m), and Table 3 (Taichi). On all datasets, PPL significantly outperforms all baselines across all image resolutions. Among the baselines, AutoLink (He et al., 2022a) also incorporates learnable connectivity priors. However, its performance is inferior due to the absence of hierarchical memory. Notably, even when masked images are used as $I_{ref}$, PPL consistently outperforms AutoLink. This demonstrates that the effectiveness of PPL is primarily attributed to its pose prior, rather than relying on cross-frame reconstruction. Additionally, we note that STT (Schmidtke et al., 2021), a baseline with human-defined priors, still underperforms PPL. Consistent with HPE (Yoo & Russakovsky, 2023), this suggests that pre-defined priors are not always optimal. PPL can learn more representative priors that outperform those manually defined.
Visualization of Pose Estimation. We provide the visualization results of pose estimation on Human3.6m (Figure ALABEL:human-vis), YouTube dog videos (Figure ALABEL:dog-vis), 11k-Hands (Figure ALABEL:hand-vis), CUB-200-2011 (Figure ALABEL:bird-vis), Horse (Figure ALABEL:horse-vis), and Flowers (Figure ALABEL:flower-vis) in Appendix A7. The visualization results demonstrate that PPL can learn categorical priors and estimate poses for various object categories, without any external annotations. For example, in Row 2, Column 2 of Figure ALABEL:human-vis, PPL correctly estimates the bowing pose of a person. In Row 2, Column 5 of Figure ALABEL:dog-vis, PPL correctly estimates the pose of a dog lowering down its head. Moreover, we found the quality of estimated poses of dogs is inferior to that of humans. This is primarily due to the significant morphological differences among various dog breeds. Additionally, dogs often perform actions such as turning around, which can lead to significant changes in pose that are difficult to accurately capture by our priors in 2D space. Additional visualization results for Flowers, Hands, and Horses in Appendix A7 further demonstrate that our PPL can be applied for various categories. Notably, flowers are rigid objects with no degrees of freedom like animals. Yet, our PPL still learns meaningful categorical priors for them.
@@ 4.2 Ablation Studies
Keypoint prior Initialization Pre Pre Pre Pre Pre Pre Rand Rand Rand Rand From mem
@@
Connectivity prior Initialization Pre Pre Pre Pre Rand Rand Pre Pre Rand Rand Rand
@@
Normalized $L2$ Error 2.51 2.66 2.58 2.70 2.54 2.61 2.68 2.72 2.75 2.83 2.56
Ablation on Prior Variants. Here, we investigate how different initializations of connectivity and keypoint priors affect pose estimation and assess whether further refining these priors enhances performance. From Table 4, we obtain several key insights: (1) Models with frozen, human-defined priors (Column 4) perform worse than our PPL, indicating that PPL learns more representative priors than those predefined by humans. (2) Refining pre-defined keypoint and connectivity priors (Column 1) outperforms our default PPL, suggesting that PPL can enhance models with human-defined priors through refinement. (3) Interestingly, randomly initializing either keypoint or connectivity priors, followed by refinement during training (Columns 5-9), yields comparable performance to models with human-defined priors. This suggests that human-defined priors may not be necessary for effective pose estimation. (4) Surprisingly, freezing randomly initialized keypoint priors also results in reasonable pose estimation accuracy, though it is still lower than PPL’s default performance (Columns 7 and 9). (5) In contrast to (4), freezing random connectivity priors prevents the model from converging, implying that connectivity priors play a more critical role in guiding pose estimations than keypoint priors.
Ablation on Number & Dimension of Vectors in Each Memory Bank. In our hierarchical memory, we used 34 memory banks. Here, we analyze the impact of the number of vectors per memory bank and the dimension of each vector on PPL’s pose estimation performance. From Figure 5 in Appendix A3, we observed that PPL remains robust across different vector counts and dimensions, although performance slightly improves with more vectors of higher dimensions. As a result, we fixed 16 vectors per memory bank, each with a dimension of 512, for all experiments.
Ablation on Number of Keypoints. We varied the number of keypoints in the pose priors from 4 to 32. The results in Figure 5 in Appendix A3 show that pose estimation accuracy improves as the number of keypoints in the prior increases. However, using 32 keypoints offers limited improvement compared to PPL with 16 keypoints.
@@ 4.3 Pose Estimation in Occluded Scenes
Visualization of Pose Estimation with Occlusion. We present the estimated poses by our PPL for occluded images as a function of the number of inference iterations in Figure LABEL:fig:occlusion-vis. Across both RandomMasking and CenterMasking, with our iterative inference strategy, PPL successfully reconstructs the occluded image parts after three iterations and meanwhile, predicts reasonable full-body poses.
@@ 5 Discussion
6 Impact Statement
We present the challenge of unsupervised learning of categorical priors for pose estimation and propose Pose Prior Learner (PPL) to address it. By leveraging learned priors, PPL enhances robustness and generalizability during inference, promoting the safe deployment of AI technologies. However, privacy concerns surrounding images and videos in these datasets, especially for human pose estimation, must be carefully addressed. Protecting personal and sensitive information remains paramount as AI advances.
@@ 6 Impact Statement
We use the Adam optimizer with a learning rate of $10^{-3}$ and a batch size of 64, training for 50 epochs. Unless specified, all images are resized to $256\times 256$. The learning rate for link weights is scaled by 512 to address the small gradients of SoftPlus near zero. We conduct experiments using the link thickness $\sigma^{2}=5\times 10^{-4}$ across all benchmark datasets, where we adopt the same definition of $\sigma$ used in (He et al., 2022a). For the hierarchical memory, we use $34$ memory banks, each of which contains $16$ vectors of dimension $512$, for all experiments.
@@
Human3.6m (Ionescu et al., 2013) is a standard benchmark dataset for human pose estimation, consisting of 3.6 million video frames. These frames include both 3D and 2D keypoints and were captured in a controlled studio environment with a static background, featuring various actors. We adhere to the approach outlined in (Zhang et al., 2018; He et al., 2022a), focusing on six activities: direction, discussion, posing, waiting, greeting, and walking. For training, we use subjects 1, 5, 6, 7, 8, and 9, and for testing, we use subject 11.
@@
Taichi (Siarohin et al., 2019b, a) consists of 3,049 training videos and 285 test videos featuring individuals performing Tai-Chi, with diverse foreground and background appearances. Following the approach in (Siarohin et al., 2021), we use 5,000 frames for training and 300 frames for testing.
@@
YouTube dog videos are videos with green backgrounds collected from YouTube to further qualitatively demonstrate the performance of PPL. Existing dog datasets are not suitable, as they often contain multiple, partially occluded dog instances. Therefore, we curated a custom dataset using 20 YouTube videos, extracting 2,000 frames for training. This dataset allows us to demonstrate PPL’s ability to learn pose priors for non-human categories without using ground truth poses. All images are trained and tested at a resolution of $256\times 256$. We use $10$ keypoints for this category and provide the visualization of the estimated poses on the test videos. The dataset will be publicly released together with other data, models, and source code.
CUB-200-2011 (Wah et al., 2011) contains 11788 images of birds. We crop and align the images according to (He et al., 2022a). We use the train/val/test split of (Choudhury et al., 2021). All images are trained and tested at a resolution of $128\times 128$.
@@
11k-Hands (Afifi, 2019), Horse (Zhu et al., 2017), and Flowers (Nilsback & Zisserman, 2008) are used for qualitative visualization. Images with multiple horses are removed and all horses are aligned to face left. All images are trained and tested at a resolution of $128\times 128$.
Appendix A3 Ablation on Memory Bank Vectors and Keypoint Numbers

### v3

@@ 3.4 Training and Inference
Our PPL is trained to jointly minimize all the four losses: the image reconstruction loss $L_{ir}$, the boundary loss $L_{b}$, the link regularization loss $L_{l}$, and the keypoint configuration reconstruction loss $L_{kr}$. We describe these losses in detail below and present their ablation results in Table A3 in Appendix A8. Additionally, we use three training techniques for stable convergence of the network. We present the details of these training techniques in Appendix A2.
Image Reconstruction Loss. If PPL correctly estimates the pose of image $I$, the reconstructed image $I_{recon}$, based on the estimated pose, should be identical to $I$. Therefore, ensuring the quality of $I_{recon}$ encourages PPL to improve its pose estimation accuracy. To achieve this, we apply a perceptual loss on the embeddings of $I$ and $I_{recon}$, extracted using a frozen feature extractor $\psi(\cdot)$ from the VGG19 network pre-trained on ImageNet Russakovsky et al. (2015). The perceptual loss is defined as: $L_{ir}=\|\psi(I_{recon})-\psi(I)\|_{1}$.
Boundary Loss. To prevent the keypoint prior from being transformed outside the image boundary, we limit the transformed keypoints to be within the image: $L_{b}=\sum_{*\in\{x,y\}}(|P^{\prime}_{i,*}|\text{ if }|P^{\prime}_{i,*}|>1,\text{ else }0)$, where $P^{\prime}_{i,x}$ and $P^{\prime}_{i,y}$ are the normalized $x$ and $y$ coordinate of the keypoint $P^{\prime}_{i}$ respectively.
Link Regularization Loss. A person’s arm always maintains a fixed length regardless of the poses. Thus, we propose the constraint that links should be assigned a high weight if they do not vary significantly in length before and after the transformation. Since limb rigidity is inherently defined in 3D space, we introduce it as an auxiliary constraint in 2D that approximates this property, primarily to accelerate convergence and stabilize training. The loss $L_{l}$ encourages the preservation of the link lengths during pose estimation. It is defined as: $L_{l}=\sum_{i,j}w_{i,j}\|l(P_{i},P_{j})-l(P^{\prime}_{i},P^{\prime}_{j})\|_{1}$, where $l(\cdot)$ is the $L2$ distance between two keypoints before and after the transformation.
@@
Iterative Inference. The memory in PPL stores learned prototypical poses, making it well-suited for correcting inaccurate pose estimates—especially under occlusion—through an iterative autoregressive process during inference (Figure 3). Iterative inference leverages this memory to progressively refine predictions alongside image reconstruction. In every iteration, we take the reconstructed image $I_{recon}$ from the last iteration (the original image $I$ for iteration 0) as the input. We infer its keypoints $T^{\prime}$ as the output keypoints of the current iteration. The hierarchical memory $M$ is used to reconstruct $T^{\prime}$ and the reconstructed keypoints $T^{\prime}_{recon}$ are used to obtain the reconstructed image $I_{recon}$. $I_{recon}$ is then used as the input for the next iteration. We keep the original occluded image $I$ as the reference image for all the iterations. $4$ iterations are used for every experiment.
In our experiments (see Section 4.3), we demonstrate that iterative inference outperforms one-step correction. This advantage stems from several key factors: (1) A single correction often fails to capture complex dependencies and may introduce additional errors. In contrast, autoregressive inference progressively refines predictions, enhancing both accuracy and consistency. (2) For tasks like pose estimation, initial predictions are often ambiguous. Iterative updates enable the model to integrate contextual cues at each step, thereby resolving uncertainties more effectively. (3) Iterative inference also encourages the model to reason about structured dependencies across spatial and temporal domains, improving robustness to scene variation, object occlusion, and background noise.
Figure 3: Overview of the iterative inference strategy in our PPL (Section 3.4). During inference, we iteratively use the reconstructed image $I_{recon}$ as input to estimate the pose $T^{\prime}$. The hierarchical memory $M$ refines the estimated pose $T^{\prime}$ and outputs $T^{\prime}_{recon}$. The original image II is used as the reference image to reconstruct the image $I_{recon}$. It is then used as the input image in the next iteration.
Training and Implementation Details. We use the Adam optimizer with a learning rate of $10^{-3}$, training for 50 epochs. The learning rate for link weights is scaled by 512 to address the small gradients of SoftPlus near zero. We conduct experiments using the link thickness $\sigma^{2}=5\times 10^{-4}$ across all benchmark datasets, where we adopt the same definition of $\sigma$ used in He et al. (2022a). For the hierarchical memory, we use $34$ memory banks, each of which contains $16$ vectors of dimension $512$. All results are averaged over 3 experiment runs. We train and test PPL on a single NVIDIA RTX A6000 GPU with the batch size of 64. Compute efficiency is reported in Table A1. See Appendix A1 for more details.
@@ 4 Experiments
For quantitative experiments, we use three public datasets: Human3.6m Ionescu et al. (2013), Taichi Siarohin et al. (2019b; a), and CUB-200-2011 Wah et al. (2011). We set the number of keypoints $N$ to be the same as He et al. (2022a). We use PPL trained on Human3.6m for all ablation studies and experiments in occluded scenes. For qualitative visualizations of the learned pose priors, we additionally include Youtube dog videos, Flowers Nilsback & Zisserman (2008), 11k-Hands Afifi (2019), and Horses Zhu et al. (2017). To train PPL, we utilize video frames as $I_{ref}$ on the Human3.6m, Taichi, and YouTube dog videos, while randomly masked images are used as $I_{ref}$ for other datasets. See Appendix A3 for dataset details.
@@ 4.1 Unsupervised Categorical Pose Estimation
We compare the keypoint detection results of PPL with other unsupervised pose estimation methods and present the results in Table 1. On all datasets, PPL outperforms all baselines across all image resolutions. We note that STT Schmidtke et al. (2021), a baseline with human-defined priors, underperforms PPL. Consistent with HPE Yoo & Russakovsky (2023), this suggests that pre-defined priors are not always optimal. PPL can learn more representative priors that outperform those manually defined. Among the baselines, AutoLink He et al. (2022a) also incorporates learnable connectivity priors; but its performance is inferior to PPL due to the absence of hierarchical memory.
To demonstrate that PPL’s advantage comes from the learned priors rather than temporal consistency in video datasets, we treat all video frames as independent images and use randomly masked images as $I_{ref}$ to train PPL on the video datasets. We denote the resulting model as PPL*. Notably the results of PPL* on Human3.6m and Taichi in Table A4 in Appendix A8 show that even without temporal consistency in video reference frames as supervision, PPL* continues to outperform AutoLink. This highlights that the effectiveness of PPL stems primarily from its pose prior, rather than from cross-frame reconstruction.
Additionally, we provide the result comparisons between our PPL and Hedlin et al. (2024) in Table A2 in Appendix A4. It is noteworthy that PPL performs competitively with Hedlin et al. (2024), which is based on priors from pre-trained text-image stable diffusion models, despite PPL being much smaller in size and relying solely on the visual modality rather than both text and vision.
We also present visual comparisons of the estimated poses between our PPL and AutoLink in Figure LABEL:fig:PPL_AutoLink. From scenes without complex backgrounds (Row 1), both AutoLink and PPL can predict correct poses. However, in more challenging settings—such as those with high-contrast backgrounds (Row 2) or where arms blend into the background (Row 3)—AutoLink occasionally misplaces keypoints on the background, resulting in implausible poses. In contrast, PPL consistently attends to the human body, leveraging its learned pose prior to enforce realistic body constraints.
Visualization of Pose Estimation. We provide the visualization results of pose estimation on Human3.6m (Figure ALABEL:human-vis), YouTube dog videos (Figure ALABEL:dog-vis), 11k-Hands (Figure ALABEL:hand-vis), CUB-200-2011 (Figure ALABEL:bird-vis), Horse (Figure ALABEL:horse-vis), and Flowers (Figure ALABEL:flower-vis) in Figure 7(a) in Appendix A5. The visualization results demonstrate that PPL can learn categorical priors and estimate poses for various object categories, without any external annotations. For example, in Row 2, Column 2 of Figure ALABEL:human-vis, PPL correctly estimates the bowing pose of a person. In Row 2, Column 5 of Figure ALABEL:dog-vis, PPL correctly estimates the pose of a dog lowering down its head. Moreover, we found that the quality of estimated dog poses is lower than that of humans, primarily because dogs exhibit much greater variability in pose, breed, size, and fur texture. Additional visualization results for Flowers, Hands, and Horses in Figure 7(a) in Appendix A5 further demonstrate that our PPL can be applied for various categories. Notably, flowers are rigid objects with no degrees of freedom like animals. Yet, our PPL still learns meaningful categorical priors for them.
@@ 4.2 Ablation Studies
Ablation on Prior Variants. We investigate how different initializations of connectivity and keypoint priors affect pose estimation and whether further refining these priors enhances performance. From Table 2, we obtain several key insights: (1) Models with frozen, human-defined priors (Column 4) perform worse than PPL, indicating that PPL learns more representative priors than those predefined by humans. (2) Refining pre-defined keypoint and connectivity priors (Column 1) outperforms our default PPL, suggesting that PPL can enhance models with human-defined priors through refinement. (3) Interestingly, randomly initializing either keypoint or connectivity priors, followed by refinement during training (Columns 5-9), yields comparable performance to models with human-defined priors. This suggests that human-defined priors may not be necessary for effective pose estimation. (4) Surprisingly, freezing randomly initialized keypoint priors also results in reasonable pose estimation accuracy, though it is still lower than PPL’s default performance (Columns 7 and 9). (5) In contrast to (4), freezing random connectivity priors prevents the model from converging, implying that connectivity priors play a more critical role in guiding pose estimations than keypoint priors.
@@
Keypoint
prior
Initialization Pre Pre Pre Pre Pre Pre Rand Rand Rand Rand From mem
@@
Connectivity
prior
Initialization Pre Pre Pre Pre Rand Rand Pre Pre Rand Rand Rand
@@
Normalized
$L2$ Error
2.51 2.66 2.58 2.70 2.54 2.61 2.68 2.72 2.75 2.83 2.56
Ablation on Number & Dimension of Vectors in Each Memory Bank. In our hierarchical memory, we used 34 memory banks. Here, we analyze the impact of the number of vectors per memory bank and the dimension of each vector on PPL’s pose estimation performance. From Figure A10 in Appendix A8, we observed that PPL remains robust across different vector counts and dimensions, although performance slightly improves with more vectors of higher dimensions. As a result, we fixed 16 vectors per memory bank, each with a dimension of 512, for all experiments.
Ablation on Number of Keypoints. We varied the number of keypoints in the pose priors from 4 to 32. The results in Figure A10 in Appendix A8 show that pose estimation accuracy improves as the number of keypoints in the prior increases. However, using 32 keypoints offers limited improvement compared to PPL with 16 keypoints.
@@ 4.3 Pose Estimation in Occluded Scenes
(a) Visualization results of poses estimation on Human3.6m. (a) Pose estimation on occluded images in Human3.6m. The first column shows the original image and its estimated pose by PPL. Columns 2-5 show the iterative inference process where the reconstructed images by PPL (Row 1 and 3) are fed back to itself for estimating poses (Rows 2 and 4) on occluded images either using CenterMasking (Row 1 and 2) or RandomMasking (Row 3 and 4). (b) The pose prior evolves as a function of training epochs (from top to bottom). (c) Comparison between PPL and AutoLink He et al. (2022a) in estimated poses on example images from Human3.6m. The left column shows testing images, the middle column and the right column show estimated poses by AutoLink and PPL respectively. All results are obtained with 16 keypoints.
@@
Visualization of Pose Estimation with Occlusion. We present the estimated poses by our PPL for occluded images as a function of the number of inference iterations in Figure LABEL:fig:occlusion-vis. Across both RandomMasking and CenterMasking, with our iterative inference strategy, PPL successfully reconstructs the occluded image parts after three iterations and meanwhile, predicts reasonable full-body poses. We include additional qualitative samples on occlusions in Figure A1 in Appendix A5, showing that PPL consistently outperforms AutoLink by reconstructing plausible poses even with partial visual information. Generally, incorrect predictions are corrected by the memory, particularly under small-area occlusions. However, severe occlusions can lead to plausible yet incorrect reconstructions, potentially compounding errors. A failure example is provided in Figure A1 in Appendix A5 where PPL fails to reconstruct the ground truth pose with large occlusion.
@@ 5 Discussion
We use the Adam optimizer with a learning rate of $10^{-3}$ and a batch size of 64, training for 50 epochs. Unless specified, all images are resized to $256\times 256$. The learning rate for link weights is scaled by 512 to address the small gradients of SoftPlus near zero. We conduct experiments using the link thickness $\sigma^{2}=5\times 10^{-4}$ across all benchmark datasets, where we adopt the same definition of $\sigma$ used in He et al. (2022a). For the hierarchical memory, we use $34$ memory banks, each of which contains $16$ vectors of dimension $512$, for all experiments. We utilize video frames as $I_{ref}$ for the Human3.6m, Taichi, and YouTube dog videos, while randomly masked images are used as $I_{ref}$ for other datasets. On Human3.6m and CUB-200-2011, we report the results in the mean $L2$ error between the predicted keypoints and the ground truth, normalized by the image size. For Taichi He et al. (2022a); Siarohin et al. (2021); He et al. (2021; 2022b); Zhang et al. (2018), we use the summed $L2$ error computed at a resolution of $256\times 256$. We excluded the DeepFashion Liu et al. (2016), CelebA Liu et al. (2015), and Zebra Zhu et al. (2017) datasets used by AutoLink because they either lack structural or background complexity, offer limited pose variation, or are redundant with existing categories. All results are averaged over 3 experiment runs. We train and test PPL on a single NVIDIA RTX A6000 GPU with the batch size of 64. The training and inference efficiency are presented in Table A1.
Table A1: Training and inference efficiency of PPL on a single NVIDIA RTX A6000 GPU with batch size of 64.
Resolution Training Time Inference Speed Inference Speed GPU Memory
(per batch) (0 iteration, img/s) (3 iterations, img/s) (usage)
$\times128$ $\sim$1.3 sec $\sim$145 $\sim$64 $\sim$9.2 GB
$\times256$ $\sim$4.0 sec $\sim$67 $\sim$24 $\sim$37.0 GB
Appendix A2 Training Techniques of our Pose Prior Learner
To ensure convergence and stability during the training of PPL, we introduce three gradient dettachment techniques: (1) To address the broken gradient issue during the quantization step from $G$ to $G^{\prime}$, we adopt the approach from VQ-VAE Van Den Oord et al. (2017). Specifically, our PPL copies the gradients of $G^{\prime}$ to $G$ for backward propagation, allowing the gradients to flow through the quantization step. (2) The hierarchical memory $M$ is updated using an exponential moving average to smooth the gradient updates, particularly during the early stages of training when $G$ can be quite noisy. This approach helps stabilize the learning process and ensures that $M$ retains more reliable information over time. (3) For $M$ to learn effective representations of $G$, it requires an accurate estimation of $T^{\prime}$, which depends on a good prior $V$ distilled from $M$. This creates a chicken-and-egg problem that complicates training. To address this, we introduce two gradient detachments to separate the training processes. First, we detach the gradients from $T$ and train the keypoint transformation and image reconstruction pathway, as shown by the red arrows in Figure 2. Second, we detach the gradients from $T^{\prime}$ to train the memory encoder and decoder, $MIX_{\text{enc}}$ and $MIX_{\text{dec}}$, as indicated by the green arrows in Figure ALABEL:fig:mem_stru in Appendix A7.
@@
Human3.6m Ionescu et al. (2013) is a standard benchmark dataset for human pose estimation, consisting of 3.6 million video frames. These frames include both 3D and 2D keypoints and were captured in a controlled studio environment with a static background, featuring various actors. We adhere to the approach outlined in Zhang et al. (2018); He et al. (2022a), focusing on six activities: direction, discussion, posing, waiting, greeting, and walking. We use subjects 1, 5, 6, 7, 8, and 9 for training; and we use subject 11 for testing.
@@
Taichi Siarohin et al. (2019b; a) consists of 3,049 training videos and 285 test videos performing Tai-Chi, with diverse foreground and background appearances. Following the approach in Siarohin et al. (2021), we use 5,000 frames for training a linear regression and 300 frames for testing.
@@
YouTube dog videos are videos with green backgrounds collected from YouTube to further qualitatively demonstrate the performance of PPL. Existing dog datasets are not suitable, as they often contain multiple, partially occluded dog instances. Therefore, we curated a custom dataset using 20 YouTube videos with a sampling frequency of 30 frames per second. We manually curate 2,000 training frames by selecting those frames with clear dog poses and removing redundant or similar ones. All videos in our custom dataset are publicly available, and are used strictly for research purposes in compliance with applicable legal and ethical guidelines. This dataset allows us to demonstrate PPL’s ability to learn pose priors for non-human categories without using ground truth poses. All images are trained and tested at a resolution of $256\times 256$. We use $10$ keypoints for this category and provide the visualization of the estimated poses on the test videos. The dataset will be publicly released together with other data, models, and source code.
CUB-200-2011 Wah et al. (2011) contains 11788 images of birds across 200 bird species with detailed annotations such as bounding boxes, part locations, and attributes. We crop and align the images according to He et al. (2022a). We use the train/val/test split of Choudhury et al. (2021). The dataset supports both aligned and non-aligned evaluation protocols, where aligned evaluation uses cropped bird regions while non-aligned evaluation relies on the original full images, making the task more challenging. The aligned subset CUB-Aligned and non-aligned subsets CUB-001/002/003/all are used for evaluation. All images are trained and tested at a resolution of $128\times 128$.
@@
11k-Hands Afifi (2019), Horse Zhu et al. (2017), and Flowers Nilsback & Zisserman (2008) are used for qualitative visualization. Images with multiple horses are removed and all horses are aligned to face left. All images are trained and tested at a resolution of $128\times 128$.
Appendix A4 Additional Comparison with Methods Using Multi-modal Knowledge
To provide broader context, we further compare our method with Hedlin et al. (2024). Hedlin et al. (2024) requires a Conditional Stable Diffusion Model pre-trained with large-scale multimodal data (image and text) to start with, thereby leveraging knowledge from an additional modality (text). The comparison results are presented in Table A2. Compared with Hedlin et al. (2024), our PPL method outperforms this baseline on Human3.6m and achieves comparable results on Tai-Chi for human pose estimation. For bird pose estimation, PPL achieves stronger performance on the CUB aligned evaluation set but lags behind on the non-aligned set. Overall, PPL employs a substantially smaller model, yet remains competitive with multimodal methods trained with extensive language supervision.
Table A2: Comparison with Hedlin et al. (2024) on keypoint detection. For Human3.6m and Taichi, we report the mean L2 error normalized by the image resolution of 128×128 and the summed L2 error at a resolution of 256×256, respectively. For CUB-200-2011, we report the mean L2 error normalized by the resolution of 128×128. We use different numbers of keypoints for the subsets of CUB-200-2011: 10 for CUB-Aligned and 4 for CUB-001, CUB-002, CUB-003, and CUB-All. The best results are in bold.
Human3.6m Taichi CUB-Aligned CUB-001 CUB-002 CUB-003 CUB-all
Hedlin et al. (2024) 4.45 234.89 5.06 10.5 11.1 10.3 5.4
PPL (ours) 1.92 293.35 3.19 19.3 18.6 17.3 10.5
Appendix A5 Additional Visualization Results of Poses under Occlusion and Priors from Other Categories
Figure A0: Comparison between PPL and AutoLink pose estimation results on occluded images on Human3.6m. The left column shows testing images, the mid column shows results of AutoLink, and the right column shows results from PPL. All results are obtained with 16 keypoints.
Figure A1: One example where PPL fails to estimate the right pose when there is a large occlusion. From left to right, we show the original unoccluded image, the largely occluded image, the result without iterative inference (iteration 0), and the final result with iterative inference (iteration 3).
(a) Additional visualization on (a) Human3.6m, (b) Youtube dog videos, (c) 11k-Hands, (d) CUB-200-211, (e) Horse, and (f) Flowers. Columns 1-4 of every subfigure show the original images and corresponding pose estimation results by PPL. The column 5 of every subfigure shows the learned prior for the category.
Appendix A6 PPL Results of Keypoint Detection on Occluded Images
Figure A7: PPL results of keypoint detection as a function of number of inference iterations on images with RandomMasking from Human3.6m. The “mask ratio" in the legend specifies the masked proportion on the $32\times 32=1024$ image patches.
Figure A8: PPL results of keypoint detection as a function of number of inference iterations on images with CenterMasking from Human3.6m. The “mask size" in the legend refers to the width and height of the masked region, on $1024$ image patches.
(a) Retrieval and distillation of the proposed hierarchical memory in our PPL. (a) The hierarchical memory $M$ is trained to reconstruct the keypoints $^{\prime}_{recon}. TT^{\prime}$ is encoded into $m$ tokens by the MLP-Mixer blocks $MIX_{enc}$. Each token $g_{i}$ retrieves its closest vector $g^{\prime}_{i}$ in memory bank $b_{i}$. The resulting $m$ vectors are decoded by the MLP-Mixer $MIX_{dec}$ into the reconstructed keypoints $T^{\prime}_{recon}$. The green arrows indicate the gradient flows during backpropagation based on the reconstruction of keypoint configurations. See Section 3.4 for training details. (b) The hierarchical memory $M$ is distilled into the keypoint prior $T$. Vectors in every memory bank $b_{i}$ are mean-pooled into one vector, and the resulting $m$ vectors are decoded by $MIX_{dec}$ into the keypoint prior $T$. See Section 3.3 for details.
Appendix A8 Additional ablation results
Number of memory bank vectors and keypoints in Human3.6m. We vary the dimensionality of the vectors in each memory bank to examine its impact on pose estimation accuracy. In addition, we analyze how the number of keypoint priors influences performance. Results are shown in Figure A10.